.Find the minimum value of with the constraints `xy+yz+zx=3

1
Expert's answer
2022-03-20T06:44:05-0400

"Let F=(x^2+y^2+z^2) +\lambda(xy+yz+zx)=0"

"F_x=0, F_y=0, F_z=0"

"2x+\lambda(y+z)=0,2y+\lambda(x+z)=0,
2z+\lambda(x+y)=0"

"\lambda=\frac{-2x}{y+z},\lambda=\frac{-2y}{x+z},\lambda=\frac{-2z}{x+y}"

Equating the equations ,we have,

"\frac{-2x}{y+z},=\frac{-2y}{x+z},=\frac{-2z}{x+y}"

Taking,

"\frac{x}{y+z},=\frac{y}{x+z}"

We have, "(x-y)(x+y+z)=0"

"x=y or x+y+z=0"

we must have x=y since x+y+z cannot be 0

Suppose x+y+z=0, squaring this we have,

"(x^2+y^2+z^2)+2(xy+yz+zx)=0"